OpenStudy (anonymous):

Which equations model this problem? A recipe that makes one pot pie calls for 0.125 g of pepper. How many pot pies can Michael make if he uses 1 g of pepper? Use p for the number of pot pies. Choose all answers that are correct. A. 1/0.125 =p B. 0.125*1=p C 0.125p=1 D. 1/p=1/0.125

OpenStudy (anonymous):

CHOOSE ALL CORRECT ANSWERS

OpenStudy (haichi):

A,D

OpenStudy (triciaal):

1 p = 0.125 grams divide each side by 0.125 1/0.125 p = 0.125/0.125 grams 1/0.125 p = 1 gram divide each side by p 1 p/p = 0.125/p 1 = 0.125 /p

OpenStudy (anonymous):

So what would the answer be?

OpenStudy (anonymous):

a d